an investor invested a total of 2700 dollars in two mutual funds. one funded 5% and the other at 2% profit.if the investors total profit was $93 how much was invested in each mutual fund
Answer by mananth(16946)

Part I 2.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
2700
Interest----- 93.00

Part I 2.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 2700 -------------1
Interest on both investments
2.00% x + 5.00% y= 93
Multiply by 100
2 x + 5 y= 9300.00 --------2
Multiply (1) by -2
we get
-2 x -2 y= -5400.00
Add this to (2)
0 x 3 y= 3900
divide by 3
y = 1300
Part I 2.00% $ 1400
Part II 5.00% $ 1300

CHECK
1400 --------- 2.00% ------- 28.00
1300 ------------- 5.00% ------- 65.00
Total -------------------- 93.00

m.ananth@hotmail.ca